Multiply
(x/3 + 3/4(3x/4 - 3/5)

Answer by fractalier(6550)

We will just FOIL it out and combine and reduce where we can...we get
(x/3 + 3/4)(3x/4 - 3/5) =
3x^2/12 - 3x/15 + 9x/16 - 9/20 =
(x^2)/4 - x/5 + 9x/16 - 9/20 =
(x^2)/4 - 16x/80 + 45x/80 - 9/20 =
(x^2)/4 + 29x/80 - 9/20